# CorMulT: A Semi-supervised Modality Correlation-aware Multimodal Transformer for Sentiment Analysis

### v1

@@ 3.2.2. Sample Generation for Contrastive Learning
3.2.3. Modality Correlation Evaluation
The unimodal representations are then fed into modality correlation evaluation (MCE) module to evaluate the correlations between different modalities. Fig. 3 shows the structure of MCE module. which consists of modality transformation layer, cross-attention layer, feedforward layers, similarity analysis layer, and loss calculation layer. The modality transformation layer aims to transform the input data into higher-level representations with dimensions equal to the hidden layer dimension. The cross-attention layer computes the attention scores across modalities to reveal the interrelations among text, audio, and images. The feedforward layers linearly transform the attention results, i.e., mapping them to the pre-defined output dimension. The goal of MCE is to minimize the triplet loss function which calculates the disparity between the predicted outputs and labels.
Figure 3. The structure of Modality Correlation Evaluation (MCE) module.
@@ 3.2.3. Modality Correlation Evaluation
Transformer Encoding.The encoder consists for three layers, and takes as inputs the $F_{A}^{\prime},F_{T}^{\prime},$ and $F_{V}^{\prime}$ to generate compact and consistent representation for each modality. The implementation details of the encoder are discussed as follows.
@@
Similarity Analysis. After the features are encoded, linear transformation is implemented to project $X_{out}$ for each modality to a shared correlation space $R^{b\times t\times o}$ for similarity comparison. Thus, we obtain $F_{A}^{\prime\prime},F_{V}^{\prime\prime},$ and $F_{T}^{\prime\prime}$ for audio, vision, and text modalities, respectively. In this space, the features from different modalities help determine the inter-modal correlations. For audio, vision and text, the mapped feature matrices in the space of $R^{b\times t\times o}$ , where $t$ is the temporal output mapping parameter and $o$ is the feature dimension output mapping parameter. $t$ adjusts the temporal dimension mapping of features, while $o$ defines the dimensionality of features in the output space. Together, these parameters enable the mapping of outputs from different modalities into a unified feature space, facilitating operations such as similarity comparison.
Then, we use cosine similarity to measure the correlations between different modalities, and obtain the pair-wised cross-modal correlation coefficients, i.e., ${Cor}_{L,A}$, ${Cor}_{L,V}$ and ${Cor}_{V,A}$.
3.2.4. Loss calculation.
@@
Concretely, the TL function measures and optimizes the relationship between the feature matrix of each modality ($F^{\prime}$) against positive ($F_{p}$) and negative ($F_{n}$) feature matrices of another modality, where ”positive” refers to the highly correlated modalities, and ”negative” indicates the uncorrelated ones, aiming to foster closer feature vectors among related modalities while enlarging the distance between unrelated ones.
For three modalities, the corresponding losses ($LOSS_{A}$, $LOSS_{T}$, $LOSS_{V}$) are derived from their similarity with the features of the other two modalities, i.e.,
(8) $\begin{split}LOSS_{A}&=\text{TL}(F_{A}^{\prime},F_{T_{p}},F_{T_{n}})+\text{TL}(F_{A}^{\prime},F_{V_{p}},F_{V_{n}})\\ LOSS_{T}&=\text{TL}(F_{T}^{\prime},F_{A_{p}},F_{A_{n}})+\text{TL}(F_{T}^{\prime},F_{V_{p}},F_{V_{n}})\\ LOSS_{V}&=\text{TL}(F_{V}^{\prime},F_{T_{p}},F_{T_{n}})+\text{TL}(F_{V}^{\prime},F_{A_{p}},F_{A_{n}})\\ \end{split}$
The final loss for modality correlation evaluation is $LOSS=\frac{LOSS_{A}+LOSS_{T}+LOSS_{V}}{3}$ which could enhance the interactions among different modalities and promote an equitable contribution from each modality towards the overall learning objective.
3.3. Correlation-aware Multimodal Transformer
@@
3.3.1. Multimodal Feature Learning
@@
$F_{m}^{\prime}=\text{Conv1D}(F_{m})$
@@
$h_{T\to V}=\text{CrossmodalTransformer}_{T\to V}(F_{T}^{\prime},F_{V}^{\prime})$
$h_{T\to A}=\text{CrossmodalTransformer}_{T\to A}(F_{T}^{\prime},F_{A}^{\prime})$
$h_{V\to T}=\text{CrossmodalTransformer}_{V\to T}(F_{V}^{\prime},F_{T}^{\prime})$
$h_{V\to A}=\text{CrossmodalTransformer}_{V\to A}(F_{V}^{\prime},F_{A}^{\prime})$
$h_{A\to T}=\text{CrossmodalTransformer}_{A\to T}(F_{A}^{\prime},F_{T}^{\prime})$
$h_{A\to V}=\text{CrossmodalTransformer}_{A\to V}(F_{A}^{\prime},F_{V}^{\prime})$
@@
3.3.2. Modality Correlation-enhanced Multimodal Fusion
@@
$\text{Cor}_{T,A},\text{Cor}_{T,V},\text{Cor}_{A,V}=\text{MCE}(F_{A},F_{T},F_{V})$
@@
Then, we obtain the feature representation $h_{m_{i}\to m_{j}}$ for each pair of modality $m_{i},m_{j}\in\{T,A,V\}$ using the cross-modal attention and self-attention layers. These features are then integrated with modality correlation coefficients to produce comprehensive representations that encompass both intra-modal and inter-modal information, i.e.,
$h_{T\to A}^{\prime}=h_{T\to A}\times\text{Cor}_{T,A}$
@@
$H_{T}^{\prime}=\text{Transformer}_{T_{mem}}(\text{Concat}(h_{A\to T}^{\prime},h_{V\to T}^{\prime}))$
@@
where $H_{T}^{\prime},H_{A}^{\prime},H_{V}^{\prime}$ are the final feature representations for text, audio, and visual modalities, respectively. Then, the learned representations for three modalities are concatenated together to obtain the comprehensive representation about the multimodal data sample, i.e., $H_{concat}=\text{Concat}(H_{T}^{\prime},H_{A}^{\prime},H_{V}^{\prime})$
@@
$y=\text{Softmax}(\text{Linear}(H_{concat}))$
where $y$ is the probability distribution over sentiment classes, and the predicted sentiment class $\hat{c}$ is given by:
$\hat{c}=\arg\max_{i}(y_{i})$
where $\hat{c}$ is the index of the highest probability in the vector $y$, corresponding to the most likely sentiment class.
4. Experiments
4.1. Dataset and Settings
4.1.1. Dataset
We conduct experiments over the CMU_MOSEI dataset which is a widely used benchmark for multimodal sentiment analysis. This dataset contains 23,453 distinct video clips that cover diverse domains, including reviews, debates and consults. Each video clip has three modalities, i.e., video, audio and text, and is labelled with seven classes of fundamental emotions with an intensity scale from -3 to 3. Table 1 provides the statistical details of CMU_MOSEI dataset. For our experiments, we allocate 70% of the dataset(16,417 clips) to the training set, 15% to the validation set(3,518 clips), and 15% to the test set(3,518 clips).
Table 1. The overview of CMU_MOSEI dataset.
@@ 4.1.2. Experiment Settings
  • MAE: the mean absolute error that measures the average deviation between the predicted results and true emotion values.
@@
4.2. Effects of Perturbation Strategy
@@
Each strategy is applied across the entire training dataset. Training parameters and model architectures are consistent across strategies to ensure a fair comparison. Table 2 presents the sentiment analysis results while using different data perturbation strategies. Among the four strategies, Strategy D achieves the highest values in terms of the four performance evaluation metrics, indicating the advantage of combining multiple data perturbation strategies.
Table 2. Performance of CorMulT when integrated with MCE under different data perturbation strategies.
Strategy $Acc@7$ $Acc@2$ $F1$ $MAE$ $Corr$
A 52.1 85.9 85.7 0.553 0.747
B 51.2 83.1 83.7 0.554 0.759
C 52.1 83.3 83.9 0.582 0.754
D 52.8 86.5 86.7 0.552 0.779
Average 51.1 83.3 83.5 0.551 0.758
4.3. Effects of Distance Measurement
The MCE module in CorMult model maps the learned audio, image, and text to a unified shared space termed the Correlation Joint Space. In this space, the semantic correlations between different data modalities can be more easily evaluated. We seek to identify the most appropriate distance metric to measure modality correlations by evaluating different similarity measures. Concretely, given two vectors $A$ and $B$, the metrics under consideration include Euclidean distance, Manhattan distance, Chebyshev distance, cosine similarity, and Mahalanobis distance, i.e.,
@@ 4.3. Effects of Distance Measurement
$d(A,B)=\left\|A-B\right\|_{2}=\sqrt{\sum_{i=1}^{n}(a_{i}-b_{i})^{2}}$
@@
$d(A,B)=\left\|A-B\right\|_{1}=\sum_{i=1}^{n}\left|a_{i}-b_{i}\right|$
@@
$d(A,B)=\left\|A-B\right\|_{\infty}=\max_{i=1}^{n}\left|a_{i}-b_{i}\right|$
@@
$\text{similarity}(A,B)=\frac{A\cdot B}{\left\|A\right\|\cdot\left\|B\right\|}$
@@
$d(A,B)=\sqrt{(A-B)\cdot C^{-1}\cdot(A-B)^{T}}$
@@
where $a_{i}$ and $b_{i}$ represent the $i^{th}$ components of vectors $A$ and $B$, respectively.
@@
Table 3 presents the results of CorMulT while using different distance measurement methods. According to the results, cosine similarity and Euclidean distance are the top performers. Given the overall assessment, cosine similarity emerges as the optimal distance metric for MCE, and could efficiently capture the inter-modal correlations.
@@
Table 3. Results of CorMulT while using different distance measurement methods.
Distance Acc@7 Acc@2 F1 MAE Corr
Euclidean Distance 52.6 86.3 86.4 54.8 77.4
Manhattan Distance 52.5 86.0 85.6 55.1 77.4
Chebyshev Distance 52.6 85.7 86.3 55.2 76.9
Cosine Similarity 52.8 86.5 86.7 55.2 77.9
Mahalanobis Distance 52.4 85.8 86.0 54.7 77.6
4.4. Parameters Tuning
In this experiment, we investigate two mapping parameters: temporal mapping dimension $t$ and feature mapping dimension $o$.
Initially, preset values for $t$ were set as $[2,4,\ldots,128]$ and for $o$ as $[50,100,\ldots,1000]$. This resulted in 165 potential $(t,o)$ combinations. For each combination, a unique MCE model was trained with identical feature extractors and Transformer Encoder architecture but different $t$ and $o$ for shared space mapping. Every model was trained and validated on the CMU_MOSEI dataset, ensuring consistent splits. To mitigate randomness, each $(t,o)$ was executed five times, averaging the performance. Considering the vast data size, only the initial 10% of the CMU_MOSEI dataset was used for training comparison.
Beyond the tri-modal triplet loss (LOSS) for evaluating MCE’s modal correlation learning capability, the epoch time was introduced as a measure of pre-training efficiency. Note, the epoch time here pertains only to the first 10% of the CMU_MOSEI dataset, equating to 2300 audio samples.
From the graphs, when $t$ is between 2-48, higher losses were observed, but post-48, the loss maintained between 0.16-0.18. Best results appeared within $t=48-64$, slightly increasing after 64. For $o$, results were generally better post-200, with a slight decline in performance within the 300-1000 range. The optimal range for $o$ was determined to be 200-300.
Considering both training loss and epoch time, the results suggest that while both $t$ and $o$ had incremental effects on epoch time, $o$ had a more pronounced impact. Hence, based on a combined assessment of loss and efficiency, the chosen dimensions for shared space are $(64,200)$.
4.5. Model Comparison
In this experiment, our CorMulT model is compared with leading existing multimodal sentiment analysis models. We also performed ablation studies within CorMulT to evaluate different modalities.
@@ 4.5. Model Comparison
Table 4. Comparison of CorMulT with multimodal sentiment analysis baseline models.
Model Modality Acc@7 Acc@2 F1 MAE Corr
EF-LSTM V+A+T 43.8 75.2 75.3 0.659 0.609
LMF V+A+T 48.1 82.1 82 0.623 0.677
TFN V+A+T 50.2 82.5 82.1 0.593 0.701
MFM V+A+T 50.3 82.4 82.3 0.568 0.717
CM-Bert V+T 50.5 83.6 83.5 0.646 0.747
MISA V+A+T 50.7 83.9 83.8 0.557 0.748
BBFN V+A+T 50.8 82.2 82.2 0.526 0.761
MULT V+A+T 51.1 83.3 83.5 0.551 0.758
ICCN V+A+T 51.2 82.2 82.4 0.565 0.713
MMLatch V+A+T 52.1 83.3 83.9 0.582 0.754
CorMulT V+A+T 52.8 86.5 86.7 0.552 0.779
Table 4 compares CorMulT with traditional benchmarks like EF-LSTM, LMF(Liu et al., 2018), and TFN(Zadeh et al., 2017), along with recent high-performing models such as CM-Bert and MISA(Hazarika et al., 2020), BBFN(Han et al., 2021), ICCN(ICCN). Across all models, our CorMulT exhibited superior performance.
Specifically, considering the F1 score and accuracy metrics (Acc@2, Acc@7), CorMulT distinctly outperformed other models. On Acc@7, CorMulT achieved 52.8%, surpassing the next best, MMLatch, at 52.1%. Similarly, for Acc@2, CorMulT scored 86.5%, notably better than MISA’s 83.9%, underlining its predictive accuracy.
Regarding the F1 score, CorMulT’s 86.7% markedly eclipsed others, showcasing its robustness, given F1’s representation as the harmonic mean of precision and recall. The MAE for CorMulT at 0.552 wasn’t distinctly superior but remained competitive, especially considering its pronounced advantages in other metrics.
Lastly, examining Corr (correlation), CorMulT excelled with a score of 0.779, substantially ahead of BBFN’s 0.761, underscoring its adeptness in comprehending multimodal data intricacies.
In conclusion, our CorMulT model manifests significant superiority in multimodal sentiment analysis, whether in accuracy, F1 scores, or correlation, further attesting to its efficacy and prominence in this domain.
4.6. Ablation Studies

### v2

@@ III-B2 Sample Generation for Contrastive Learning
III-B3 Modality Correlation Evaluation
The unimodal representations are then fed into modality correlation evaluation (MCE) module to evaluate the correlations between different modalities. Fig. 3 shows the structure of MCE module. which consists of modality transformation layer, cross-attention layer, feedforward layers, similarity analysis layer, and loss calculation layer. The modality transformation layer transforms the input data into higher-level representations with dimensions equal to the hidden layer dimension. The cross-attention layer calculates the attention scores across modalities to reveal the interrelations among text, audio, and images. The feedforward layers linearly transform the attention results, i.e., mapping them to the pre-defined output dimension. The goal of MCE is to minimize the triplet loss function which calculates the disparity between the predicted outputs and labels.
Figure 3: The structure of Modality Correlation Evaluation (MCE) module which learns the modality correlation and outputs the coefficients among modalities.
@@ III-B3 Modality Correlation Evaluation
Transformer encoding. The encoder consists of three layers, and takes as inputs $F_{A}^{\prime},F_{T}^{\prime},$ and $F_{V}^{\prime}$ to generate compact and consistent representation for each modality. The implementation details of the encoder are discussed as follows.
@@
Similarity analysis. After the features are encoded, linear transformation is implemented to project $X_{out}$ for each modality to a shared correlation space $R^{b\times t\times o}$ for similarity comparison, where $t$ is the temporal output mapping parameter and $o$ specifies the dimensionality of features in the output space. Thus, we obtain $F_{A}^{\prime\prime},F_{V}^{\prime\prime},$ and $F_{T}^{\prime\prime}$ for audio, vision, and text modalities, respectively, space $R^{b\times t\times o}$. Then, we use cosine similarity to measure the correlations between different modalities, and obtain the pair-wised cross-modal correlation coefficients, i.e., ${Cor}_{T,A}$, ${Cor}_{T,V}$ and ${Cor}_{V,A}$.
III-B4 Loss Calculation.
@@
Concretely, the TL function measures and optimizes the relationship between the feature matrix of each modality ($F^{\prime}$) against positive ($F_{p}$) and negative ($F_{n}$) feature matrices of another modality, where ”positive” refers to the highly correlated modalities, and ”negative” indicates the uncorrelated ones. The aim is to foster closer feature vectors among related modalities while enlarging the distance between unrelated ones.
The TL function is defined as:
$\text{TL}(F^{\prime},F_{p},F_{n})=\max(0,\text{d}(F^{\prime},F_{p})-\text{d}(F^{\prime},F_{n})+\mu)$ (8)
where d represents a distance function and margin is a predefined threshold, and $\mu$ represents the margin set as 0.2 in our experiment.
For three modalities, the corresponding losses, i.e., $LOSS_{A}$, $LOSS_{T}$, and $LOSS_{V}$, are derived from their similarity with the features of the other two modalities:
$\begin{split}LOSS_{A}&=\text{TL}(F_{A}^{\prime},F_{T_{p}},F_{T_{n}})+\text{TL}(F_{A}^{\prime},F_{V_{p}},F_{V_{n}})\\ LOSS_{T}&=\text{TL}(F_{T}^{\prime},F_{A_{p}},F_{A_{n}})+\text{TL}(F_{T}^{\prime},F_{V_{p}},F_{V_{n}})\\ LOSS_{V}&=\text{TL}(F_{V}^{\prime},F_{T_{p}},F_{T_{n}})+\text{TL}(F_{V}^{\prime},F_{A_{p}},F_{A_{n}})\\ \end{split}$ (9)
The final loss for modality correlation evaluation is $LOSS=\frac{LOSS_{A}+LOSS_{T}+LOSS_{V}}{3}$, which could enhance the interactions among different modalities and ensure an equal contribution from each modality towards the overall learning objective.
III-C Correlation-aware Multimodal Transformer
@@
III-C1 Multimodal Feature Learning
@@
$F_{m}^{\prime}=\text{Conv1D}(F_{m})$ (10)
@@
[…]
@@
III-C2 Modality Correlation-enhanced Multimodal Fusion
@@
$\text{Cor}_{T,A},\text{Cor}_{T,V},\text{Cor}_{A,V}=\text{MCE}(F_{A},F_{T},F_{V})$ (12)
@@
Then, we obtain the feature representation $h_{m_{i}\to m_{j}}$ for each pair of modality $m_{i},m_{j}\in\{T,A,V\}$ using the cross-modal attention and self-attention layers. These features are integrated with modality correlation coefficients to produce comprehensive representations that encompass both intra-modal and inter-modal information, i.e.,
$h_{T\to A}^{\prime}$ $=h_{T\to A}\times\text{Cor}_{T,A}$ (13)
@@
$H_{T}^{\prime}$ $=\text{Transformer}_{T_{mem}}(\text{Concat}(h_{A\to T}^{\prime},h_{V\to T}^{\prime}))$ (14)
@@
where $H_{T}^{\prime},H_{A}^{\prime},H_{V}^{\prime}$ are the final feature representations for text, audio and visual modalities, respectively. Then, the learned representations for three modalities are concatenated together to obtain the comprehensive representation about the multimodal data sample, i.e., $H_{concat}=\text{Concat}(H_{T}^{\prime},H_{A}^{\prime},H_{V}^{\prime})$
@@
$y=\text{Softmax}(\text{Linear}(H_{concat}))$ (15)
where $y$ is the probability distribution over sentiment classes, and the predicted sentiment class $\hat{c}$ is obtained by:
$\hat{c}=\arg\max_{i}(y_{i})$ (16)
where $\hat{c}$ is the index of the highest probability in vector $y$, corresponding to the most likely sentiment class.
IV Experiments
IV-A Dataset and Settings
IV-A1 Dataset
We conduct experiments over the CMU_MOSEI dataset which is a widely used benchmark for multimodal sentiment analysis. This dataset contains 3,228 distinct video clips that cover diverse domains, including reviews, debates and consults. Each video clip has three modalities, i.e., video, audio and text, and is labelled with seven classes of fundamental emotions with an intensity scale from -3 to 3. Within these clips, there are a total of 23,453 sentences, with each sentence along with its corresponding audio and video segments jointly constituting a sample. Table I provides the statistical details of CMU_MOSEI dataset. For our experiments, we allocate 70% of the dataset(16,417 sentences) to the training set, 15% to the validation set(3,518 sentences), and 15% to the test set(3,518 sentences).
TABLE I: The overview of CMU_MOSEI dataset.
@@ IV-A2 Experiment Settings
IV-B Effects of Perturbation Strategy
@@
Each strategy is applied across the entire training dataset. Training parameters and model architectures are consistent across strategies to ensure a fair comparison. Table II presents the sentiment analysis results while using different data perturbation strategies. Among the four strategies, Strategy D achieves the highest values in terms of all five performance evaluation metrics, indicating the advantage of combining multiple data perturbation strategies.
TABLE II: The results of CorMulT when using different data perturbation strategies for contrastive learning in MCE module.
Strategy $Acc@7$ $Acc@2$ $F1$ $Corr$
A 52.1 85.9 85.7 0.747
B 51.2 83.1 83.7 0.759
C 52.1 83.3 83.9 0.754
D 52.8 86.5 86.7 0.779
IV-C Effects of Distance Measurement
The MCE module in CorMult model maps the learned audio, image, and text to a unified shared space in which the semantic correlations between different data modalities can be well evaluated. We seek to identify the most appropriate distance metric to measure modality correlations. Concretely, given two vectors $A$ and $B$, the metrics under consideration include Euclidean distance, Manhattan distance, Chebyshev distance, cosine similarity, and Mahalanobis distance, i.e.,
@@ IV-C Effects of Distance Measurement
$d(A,B)=\left\|A-B\right\|_{2}=\sqrt{\sum_{i=1}^{n}(a_{i}-b_{i})^{2}}$ (17)
@@
$d(A,B)=\left\|A-B\right\|_{1}=\sum_{i=1}^{n}\left|a_{i}-b_{i}\right|$ (18)
@@
$d(A,B)=\left\|A-B\right\|_{\infty}=\max_{i=1}^{n}\left|a_{i}-b_{i}\right|$ (19)
@@
$\text{similarity}(A,B)=\frac{A\cdot B}{\left\|A\right\|\cdot\left\|B\right\|}$ (20)
@@
$d(A,B)=\sqrt{(A-B)\cdot C^{-1}\cdot(A-B)^{T}}$ (21)
@@
where $d(A,B)$ denotes the distance between vectors $A$ and $B$, and $a_{i}$ and $b_{i}$ represent the $i^{th}$ components of vectors $A$ and $B$, respectively.
@@
Table III presents the results of CorMulT while using different distance measurement methods. According to the results, cosine similarity emerges as the optimal distance metric for MCE, and could efficiently capture the inter-modal correlations.
@@
TABLE III: Results of CorMulT while using different distance measurement methods.
Distance Acc@7 Acc@2 F1 Corr
Euclidean Distance 52.6 86.3 86.4 77.4
Manhattan Distance 52.5 86.0 85.6 77.4
Chebyshev Distance 52.6 85.7 86.3 76.9
Cosine Similarity 52.8 86.5 86.7 77.9
Mahalanobis Distance 52.4 85.8 86.0 77.6
IV-D Parameters Tuning
Temporal mapping dimension $t$ and feature mapping dimension $o$ are two important parameters in STAD model. We thus conduct experiments to find appropriate values for them using the grid search method. Fig. 4 illustrates the training loss and epoch time(the training time for each epoch during the training of MCE) of STAD while varying $t$ from 2 to 128, and varying $o$ from 50 to 1000. To mitigate the randomness, for each combination $(t,o)$, we run experiments five times, and report the average results. Considering the large number of combinations, only the first 10% of the CMU_MOSEI dataset was used for tuning parameters, thus reducing the time cost. According to Fig. 4, when $t$ is between 2-48, higher losses were observed, but when $t$ is higher than 48, the loss maintained between 0.16-0.18. Best results appeared within $t=48-64$, slightly increasing after 64. For $o$, results were generally better when $o$ is higher than 200, with a slight decline in performance within the 300-1000 range. The optimal range for $o$ was determined to be 200-300. As the epoch time is also increasing slightly as $t$ and $o$ increase, we need to find the parameters that achieve the comparatively best results but should not introduce high latency. Hence, based on a combined assessment of loss and efficiency, the chosen dimensions for shared space are $(64,200)$.
Figure 4: The heatmap of training loss and epoch time while varying temporal mapping dimension and feature mapping dimension.
IV-E Model Comparison
CorMulT model is compared with multiple representative multimodal sentiment analysis models, including EF-LSTM[32], LMF[33], TFN[34], MFM[35], CM-Bert[36], MISA[8], BBFN[4], MULT[18], ICCN[37] and MMLatch[5]. Table IV presents the results of CorMulT and baseline methods, where V+A+T indicates that the model utilizes vision, audio and text modalities. CM-Bert only uses vision and text modalities, i.e., V+T. Among all models, CorMulT exhibits superior performance. Specifically, considering the F1 score and accuracy metrics (Acc@2 and Acc@7), CorMulT distinctly outperforms other models. For Acc@7, CorMulT achieves 52.8%, surpassing the next best, i.e., MMLatch, at 52.1%. Similarly, for Acc@2, CorMulT scores 86.5%, notably better than MISA’s 83.9%. Regarding the F1 score, CorMulT’s 86.7% markedly eclipsed others, showcasing its robustness since F1 is the harmonic mean of precision and recall. Lastly, examining Corr (correlation), CorMulT excels with a score of 0.779, substantially ahead of BBFN’s 0.761, underscoring its adeptness in comprehending multimodal data intricacies. In sum, CorMulT model manifests significant superiority in multimodal sentiment analysis in terms of accuracy, F1 scores, and correlation.
@@ IV-E Model Comparison
TABLE IV: Comparison of CorMulT with multiple baseline models for multimodal sentiment analysis.
Model Modality Acc@7 Acc@2 F1 Corr
EF-LSTM V+A+T 43.8 75.2 75.3 0.609
LMF V+A+T 48.1 82.1 82 0.677
TFN V+A+T 50.2 82.5 82.1 0.701
MFM V+A+T 50.3 82.4 82.3 0.717
CM-Bert V+T 50.5 83.6 83.5 0.747
MISA V+A+T 50.7 83.9 83.8 0.748
BBFN V+A+T 50.8 82.2 82.2 0.761
MULT V+A+T 51.1 83.3 83.5 0.758
ICCN V+A+T 51.2 82.2 82.4 0.713
MMLatch V+A+T 52.1 83.3 83.9 0.754
CorMulT V+A+T 52.8 86.5 86.7 0.779
IV-F Ablation Studies